# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v2

@@ 2.2.2 Granularity of quantization
2.3 FP4 Quantization and Variant Formats
@@
FP4 can better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, and MXFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, FP4 is more suitable than INT4 as a quantization data type.
(a) Full-precision weight
(b) INT4 quant weight
(c) MXFP4 quant weight
Figure 5: Weight distribution of a q_proj layer of LlaMA-3.2-1B. (a): Weight distribution in full-precision(BF16). (b): Weight distribution after INT4 quantization (group size=32), compared with MXFP4. (c): Weight distribution after MXFP4 quantization.
@@
3 Extensive Evaluations
Quantization settings We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
Models and tasks. We conduct experiments on LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B. We evaluate their perplexity on the WikiText-2 dataset [30] and assess their accuracy on several zero-shot evaluation tasks, including commonsense reasoning, mathematics, sentiment analysis, and question answering. Specifically, we utilize the following datasets for evaluation: MMLU [31], SST2 [32], MathQA [33], ARC‑C [34], HellaSwag [35], WinoGrande [36], BoolQ [37], LAMBADA [38], and PIQA [39].
Implementation. To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows the OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank decomposition, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
3.1 Overall performance
@@
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 8: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
@@
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 9: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank decomposition compensation. A series of experiments is conducted to evaluate these combinations. Table 8 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations. Table 9 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
@@
Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank decomposition. This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
Between the two quantization error mitigation methods, GPTQ clearly plays a more significant role than low-rank adaptation. Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings. When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model.
In addition, using both optimized rotation and scaling improves performance over using only optimized rotation. However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not perform well together with rotation in W4A4 settings.
Model precision
Wiki2
PPL($\downarrow$)
Zero-Shot Accuracy($\uparrow$)
General Sentiment Math Reasoning QA Avg.
MMLU SST2 MathQA PiQA HellaS WinoG BoolQ Lambada Arc-C
Llama-
3.2-1B
fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
w4a4 11.73 29.21 55.28 25.43 69.15 57.93 56.43 59.02 56.78 32.34 49.06
Llama-
3.2-3B
fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
w4a4 8.9 46.7 68.58 30.55 73.99 69.54 65.43 68.75 65.32 41.55 58.93
Qwen-
2.5-3B
fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
w4a4 9.22 56.32 69.15 32.5 74.16 67.52 62.12 65.41 57.77 42.83 58.64
Table 10: Comparison of symmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
In Table 10, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank decomposition.
3.2 Symmetric vs. asymmetric
Granularity Metric rotation+scaling+GPTQ Rotation+Scaling+GPTQ+Low-rank
Sym W-Asym A-Asym Asym Sym W-Asym A-Asym Asym
channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7 12.63 12.65 11.73 11.64
Arc-Challenge Score(%) 31 31.7 33.3 33.7 31.9 29.9 32.3 32.7
@@
Table 11: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [34] accuracy of LlaMA-3.2-1B
@@
In Table 11, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively. The results demonstrate that asymmetric quantization for activations leads to a significant improvement over symmetric quantization, while the performance gain for weights is relatively minor. We attribute this to the fact that activation distributions tend to be asymmetric, whereas weights are generally more symmetric in distribution.
@@
3.3 Granularity
As illustrated in Fig. 6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 12, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 6: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
@@ 3.3 Granularity
Average Extra bits
1 0.5 0.25 0.125 0.0625
Table 12: Average extra bits introduced by FP32 scaling factors of different granularities
3.4 FP4 vs. INT4
@@
First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods. Table 13 demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel). We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results. For finer granularity scenarios, FP4 and INT4 show very small differences, and both the format of the scaling factor and the size of the group significantly impact performance. Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format. With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision. By contrast, MXFP4 uses the E8M0 format, which restricts scale factors to powers of two, resulting in much coarser quantization. Moreover, NVFP4 employs a two-level micro-block scaling strategy: it first applies a per-tensor scaling factor (FP32) to compensate for the smaller range that E4M3 scales can represent. NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [40].
Quantization
error
mitigation
Data Format
INT4
channel, FP32
FP4
channel, FP32
INT4
g16, FP32
FP4
g16, FP32
MXFP4
g32, E8M0
FP4
g32, E4M3
FP4
g16, E8M0
NVFP4
g16, E4M3
RTN 256.97 135.68 11.75 11.34 15.17 12.04 14.52 11.54
GPTQ 161.04 101.36 11.47 10.87 13.24 11.36 12.77 10.98
low-rank 179.87 132.14 11.69 11.32 15.56 11.98 14.67 11.48
GPTQ+low-rank 139.31 96.33 10.92 10.76 12.69 11.18 12.25 10.83
@@
Table 13: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor. And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
Quantization
error
mitigation
Data Format
NVFP4
W:E4M3, A:E4M3
FP4
W:E2M1, A:E4M3
FP4
W:E2M1, A:E2M1
FP4
W:INT4, A:E4M3
FP4
W:INT4, A:INT4

### v3

@@ II-B2 Granularity of quantization
II-C FP4 Quantization and Variant Formats
@@
FP4 can better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, MXFP4 and NVFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, FP4 is more suitable than INT4 as a quantization data type.
((a))
((b))
((c))
((d))
Figure 5: Weight distribution of a q_proj layer of LlaMA-3.2-1B. (a): Weight distribution in full-precision(BF16). (b): Weight distribution after INT4 quantization (group size=32), compared with MXFP4. (c): Weight distribution after MXFP4 quantization. (d): Weight distribution after NVFP4 quantization.
@@
III Extensive Evaluations
Evaluation Scope. Based on the previous review and the defined experiment scope, we conducted extensive evaluations to systematically investigate the effects of Post-Training Quantization (PTQ) on large language models. Our analysis is organized into three main parts. In Section III-A, we evaluate several popular methods within the two-step quantization framework, providing a baseline for comparison. Section III-B performs an ablation study across different quantization settings, examining how these choices influence model performance. Finally, Section III-C applies the evaluated methods to FP4 and its variants, further exploring the potential and limitations of FP4 quantization.
Quantization settings. We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
Models and tasks. We conduct experiments on LLaMA-3.2 (1B, 3B), LLaMA-3.1 (8B, 70B), and Qwen2.5-3B. Additionally, we perform experiments on the MoE model Qwen1.5-MoE-A2.7B to further extend our evaluation. We evaluate their perplexity on the WikiText-2 dataset [23] and assess their accuracy on several zero-shot evaluation tasks, including commonsense reasoning, mathematics, sentiment analysis, and question answering. Specifically, we utilize the following datasets for evaluation: MMLU [12], SST2 [33], MathQA [2], ARC‑C [5], HellaSwag [40], WinoGrande [30], BoolQ [7], LAMBADA [27] and PIQA [4].
Implementation. To calibrate the GPTQ [10] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows the OSTQuant [13] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank compensation, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [38] calibration scheme with $\alpha$=0.5.
III-A Methods evaluation under the two-step quantization framework
We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank compensation. A series of experiments is conducted to evaluate these combinations. Table VIII presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations.
To assess the statistical robustness of these results, we repeated the experiments of Table VIII across five different random seeds. Fig. 7 visualizes the resulting error bars, where each bar represents the standard deviation of the Wiki2-PPL across the runs. The relatively small error bars indicate that the performance of each method is consistent and stable, confirming that the observed differences in Table VIII are statistically reliable. Table 7 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
In Table IX, we further conduct experiments on different models. The experiments of LLaMA-3.2-1B, LLaMA-3.2-3B, LLaMA-3.1-8B, LLaMA-3.1-70B and Qwen2.5-3B use the optimized rotation and scaling, and combine them with GPTQ and an additional low-rank compensation. For the MoE model Qwen1.5-MoE-A2.7B, given that the key advantage of MoE models lies in the fact that the actual computational cost scales only with the number of activated experts and performs full optimization of rotation and scaling parameters and GPTQ would be computationally expensive in quantization process compared to models with comparable computational budgets, we opt for more resource-efficient strategies which only use random rotation and low-rank compensation during PTQ.
TABLE VIII: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
@@ III-A Methods evaluation under the two-step quantization framework
PT QEM Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Figure 6: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
Figure 7: Error bars showing the statistical robustness of the methods in Table VIII, computed over five runs with different random seeds.
@@
PT QEM Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
TABLE IX: Evaluation of performance under different models.
Metric Llama- 3.2-1B Llama- 3.2-3B Qwen2.5- 3B Qwen1.5- MoE-A2.7B Llama- 3.1-8B Llama- 3.1-70B
FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4
WikiText2 PPL($\downarrow$) 9.76 11.73 7.82 8.90 8.03 9.22 7.22 8.88 6.24 7.23 2.81 4.41
Zero-Shot Accuracy($\uparrow$) General MMLU 36.76 29.21 54.06 46.70 65.13 56.32 60.86 51.38 63.32 56.98 75.26 70.57
Sentiment SST2 66.86 55.28 73.62 68.58 90.37 69.15 67.89 60.32 77.18 59.52 82.11 75.69
Math MathQA 28.94 25.43 34.47 30.55 37.15 32.50 35.01 29.21 39.50 34.91 51.66 45.36
Reasoning PiQA 74.65 69.15 77.48 73.99 78.73 74.16 80.69 76.39 81.18 78.51 84.22 81.12
HellaS 63.67 57.93 73.58 69.54 73.55 67.52 77.30 72.08 78.92 76.05 85.03 83.37
WinoG 60.38 56.43 69.46 65.43 68.75 62.12 69.38 62.59 73.4 68.9 79.79 76.24
QA BoolQ 63.67 59.02 72.97 68.75 77.58 65.41 79.60 73.98 82.05 78.75 85.5 83.39
Lambada 62.10 56.78 70.08 65.32 66.80 57.77 71.38 61.69 75.33 72.48 78.75 76.75
Arc-C 36.01 32.34 46.08 41.55 47.35 42.83 44.37 39.76 53.75 48.29 64.85 59.98
Average 54.78 49.06 63.53 58.93 67.27 58.64 65.16 58.60 69.40 63.82 76.35 72.50
In the following subsections, we provide a comprehensive and systematic analysis of the experimental results from multiple perspectives.
III-A1 Pre-quantization transformation
First, we analyze the performance of different pre-quantization transformation methods.
Under optimization, using both optimized rotation and optimized scaling yields the best performance. As shown in Table VIII, applying optimized scaling together with optimized rotation always yields better performance. This is mainly because, after optimization, scaling can effectively fine-tune the inter-channel magnitude distribution of the rotated tensor, further improving the results and achieving better spatial utilization. Using only rotation also achieves significantly better performance compared with random rotation, which is consistent with the observation in SpinQuant [22] that optimization helps eliminate the high variance introduced by random rotations.
Without optimization, it is better to use random rotation only. When the model is large, performing optimization is often impractical; therefore, we can only rely on random rotations and scaling factors obtained through calibration. From Table 7, we can see that in the absence of optimization, rotation alone performs better than rotation combined with scaling. We believe this is because when the scaling is applied after rotation, we hope it can fine-tune the rotated matrix. However, the calibration method becomes inaccurate in this case because outliers are averaged across channels, and the calculation of each channel’s scaling parameter cannot simply be based on its maximum value.
III-A2 Quantization error mitigation
Comparison between self-compensation and low-rank compensation. Between the two quantization error mitigation methods, we found that, except for the per-channel quantization case using only scaling, the effectiveness of low-rank compensation is often inferior to GPTQ. Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings. We believe the reason is that the low-rank method used in the experiments performs SVD based on the weight quantization error. It cannot accurately approximate the loss error, unlike methods like GPTQ that directly analyze the loss error. For low-rank compensation methods starting from the loss error, better results might be achieved, but it would require a relatively complex computational process, which we do not evaluate here.
Combination of self-compensation and low-rank compensation. When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model. We believe this is because low-rank compensation on the weight quantization error after GPTQ further compensates for the quantization error and leads to better results.
III-A3 Relationship between the two-step
@@
The performance contributions of these two steps are additive and mutually complementary. Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank compensation. This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
III-A4 Inference cost analysis
Table X presents the theoretical analysis of inference cost for the evaluated methods. For experiments in Table VIII and 7, we only apply online rotation to the down-projection’s input in the MLP in each block. According to SpinQuant [22], which also applies online Hadamard rotation after ROPE to enhance key-value cache quantization performance, in addition to the inputs of the down-projection. And SpinQuant [22] has tested that these two online Hadamard processes introduce only a modest $\sim$8% increase in latency.
Besides, the overhead of low-rank compensation is the most noteworthy aspect to discuss. As shown in Table X, we compute the additional memory cost and computation (FLOPs) of the low-rank computation. In our experiments, we set $r=32$. For Llama-3.2-1B, the low-rank compensation introduces approximately 1.41 M additional parameters per decoder layer, compared to about 61 M parameters in all linear layers, costing roughly 2.3% memory overhead additionally in FP16. After quantization, it costs roughly 9.2% memory overhead additionally. And it has an overhead computation ratio of $3.1\%$ in FP16.
TABLE X: Inference Cost Analysis.
Method Implementation Inference Overhead
Scaling Fused into weight None
Rotation Most: Fused into weight Fused rotation: None
Few: Online Hadamard rotation Online rotation: Fast Hadamard
GPTQ Directly weight changed None
Low-rank Additional memory: $\begin{aligned} &A\in\mathbb{R}^{c_{\text{in}}\times r}\\ &B\in\mathbb{R}^{r\times c_{\text{out}}}\end{aligned}$ Memory: $\begin{aligned} \text{ratio}=\frac{r\times(c_{\text{in}}+c_{\text{out}})}{c_{\text{in}}c_{\text{out}}}\end{aligned}$
Additional computation: XABXAB Computation: $\begin{aligned} \text{ratio}=\frac{r\times(c_{\text{in}}+c_{\text{out}})}{c_{\text{in}}\times c_{\text{out}}}\end{aligned}$
III-B Evaluation of quantization settings
III-B1 Symmetric vs. asymmetric
@@ III-B Evaluation of quantization settings
TABLE XI: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc-challenge [5] accuracy of LlaMA-3.2-1B
Granularity Metric Rotation+Scaling+GPTQ
Sym W-Asym A-Asym Asym
Channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7
Arc-C Score(%) 31 31.7 33.3 33.7
Granularity Metric Rotation+Scaling+GPTQ+Low-rank
Sym W-Asym A-Asym Asym
Channel Wiki2-PPL($\downarrow$) 12.63 12.65 11.73 11.64
Arc-C Score(%) 31.9 29.9 32.3 32.7
@@
In Table XI, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively. The results demonstrate that asymmetric quantization for activations leads to a significant improvement over symmetric quantization, while the performance gain for weights is relatively minor. We attribute this to the fact that activation distributions tend to be asymmetric, whereas weights are generally more symmetric in distribution.
@@
III-B2 Granularity
As illustrated in Fig. XII, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 8, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 8: Average extra bits introduced by FP32 scaling factors of different granularities
TABLE XII: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
@@ III-B2 Granularity
Average Extra bits 1 0.5 0.25 0.125 0.0625
III-C Extending Experiments in FP4
III-C1 FP4 vs. INT4
@@ III-C Extending Experiments in FP4
TABLE XIII: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor. And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
Data Format INT4 channel, FP32 FP4 channel, FP32 INT4 g16, FP32 FP4 g16, FP32
RTN 256.97 135.68 11.75 11.34
GPTQ 161.04 101.36 11.47 10.87
Low-rank 179.87 132.14 11.69 11.32
GPTQ+Low-rank 139.31 96.33 10.92 10.76
Data Format MXFP4 g32, E8M0 FP4 g32, E4M3 FP4 g16, E8M0 NVFP4 g16, E4M3
RTN 15.17 12.04 14.52 11.54
GPTQ 13.24 11.36 12.77 10.98
Low-rank 15.56 11.98 14.67 11.48
GPTQ+Low-rank 12.69 11.18 12.25 10.83
First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods.
FP4 performs better than INT4 without pre-quantization transformation in per-channel setting. Table XIII demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel). We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results.
For finer granularity scenarios, FP4 and INT4 show very small differences. Table XIII demonstrates that FP4 and INT4 show minimal differences at finer granularities, with FP4 performing only slightly better than INT4. We attribute this to the fact that fine-grained quantization helps mitigate the impact of outliers. Consequently, the advantage of FP4 in handling outliers is reduced under finer granularity, resulting in only a marginal performance improvement over INT4.
Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format. We believe that both the format of the scaling factor and the size of the group significantly impact performance. With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision. By contrast, MXFP4 uses the E8M0 format, which restricts scale factors to powers of two, resulting in much coarser quantization. Moreover, NVFP4 employs a two-level micro-block scaling strategy: it first applies a per-tensor scaling factor (FP32) to compensate for the smaller range that E4M3 scales can represent. NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [1].
TABLE XIV: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor.
Evaluate Methods Wiki2‑PPL($\downarrow$)
Pre-quantization transformation Quantization error mitigation INT4 channel, FP32 FP4 channel, FP32 MXFP4 g32, E8M0 NVFP4 g16, E4M3
Scaling GPTQ 134.85 89.22 13.24 10.97
Scaling Low-rank 165.4 104.55 14.91 11.36
Scaling GPTQ+Low-rank 124.56 114.96 12.58 10.81
Optimized rotation+scaling GPTQ 11.8 12.13 12.37 11.04
Optimized rotation+scaling Low-rank 14.58 15.15 15.43 12.57
Optimized rotation+scaling GPTQ+Low-rank 11.73 12.13 12.29 11.01
Then, we apply pre-quantization transformation together with quantization error mitigation on FP4 and main FP4 variants to compare with INT4, and the results are shown in Table XIV.
Rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel). As argued in AMXFP4 [16], grouping activation tensors into small micro-scaled units mitigates outliers like rotation. For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
FP4 does not show its advantage under rotation. Although MXFP4 and NVFP4 perform better than INT4 (per-channel), it is observed that under the same channel-wise setting, INT4 performs better than FP4 after rotation, while when applying only scaling, FP4 still outperforms INT4. We believe this is because rotation can largely eliminate outliers, to the extent that the advantage of FP4 over INT4 when dealing with outliers diminishes. Consequently, INT4 demonstrates stronger performance in handling more uniform distributions.
III-C2 Exploration of FP4 scaling factor
In the analysis of Table XIII in Section III-C1, we observed that the scaling factor has a significant impact, and we conduct further exploratory research on this aspect. We further reduced the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting. As shown in Table XV, applying 4-bit scaling factors to both weights and activations leads to very poor performance. When applied only to weights, the performance remains moderate but is worse compared to NVFP4.
TABLE XV: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. E4M3, E2M1, and INT4 refer to the format of the scaling factor. W and A denote weight and activation, respectively. All of these follow NVFP4’s approach, employing a two-level scaling scheme.
Quantization error mitigation Data Format
NVFP4 W:E4M3 A:E4M3 FP4 W:E2M1 A:E4M3 FP4 W:E2M1 A:E2M1 FP4 W:INT4 A:E4M3 FP4 W:INT4 A:INT4